OpenStudy (sh3lsh):

Determine the Taylor Series about x=2 for 1/(1-x) and the radius of convergence

OpenStudy (sh3lsh):

So I did the derivatives of f(x) and arrive that the general derivative pattern is n!/(1-x)^(n+1) So what's next?

OpenStudy (owen3):

use the taylor series formula $$ \sum_{n=0}^{\infty } \frac{f^{(n)}(a)(x-a)^n}{n!} $$

OpenStudy (sh3lsh):

f^(n)(a)=(-1)^(n+a)n! will continue..

OpenStudy (owen3):

by the way I got a different taylor expansion

OpenStudy (sh3lsh):

In regards to my first comment or third? My third are the coefficients, what is that called?

OpenStudy (owen3):

$$ \frac{1}{1-x }= -1+(x-2)- \left( x-2 \right) ^{2}+ \left( x-2 \right) ^{3}- \left( x-2 \right) ^{4}+ \left( x-2 \right) ^{5}+...$$

OpenStudy (sh3lsh):

So, what's the purpose of finding the general equation for the derivatives if we have to plug in 2 for each one of them? 1. find multiple derivatives 2. plug in point. 3. find general equation of that. 4. divide by n! and multiply by (x-x0)^n Is that a general case? Wouldn't that just result in -1^(n+1)*(x-2)^n ?

OpenStudy (owen3):

I agree with your general derivative $$ f(x) = \frac{1}{\left( 1-x \right) } \\ f '(x) = \frac{1}{\left( 1-x \right) ^{2}} \\ f ''(x) = \frac{2}{\left( 1-x \right) ^{3}} \\ f '''(x) = \frac{6}{\left( 1-x \right) ^{4}} \\ ... \\ \\ f^{(n)}(x) = \frac{n!}{\left( 1-x \right) ^{n+1}} $$

OpenStudy (sh3lsh):

Why does one even bother solving for the general derivative?

OpenStudy (sh3lsh):

How were you able to ascertain your third comment without fully solving the problem?

OpenStudy (owen3):

I am using a computer algebra system, like wolfram or Maple

OpenStudy (sh3lsh):

Oh.

OpenStudy (owen3):

but it can be done 'by hand', it is instructive to do a few of these problems by hand

OpenStudy (sh3lsh):

yeah, I have do this by hand..

OpenStudy (owen3):

ok so we have the general derivative, now we need to plug in x=2 and find a pattern

OpenStudy (owen3):

$$f(x) = \frac{1}{\left( 1-x \right) } \implies f(2) = -1 \\ f '(x) = \frac{1}{\left( 1-x \right) ^{2}} \implies f '(2) = 1 \\ f ''(x) = \frac{2}{\left( 1-x \right) ^{3}} \implies f ''(2) = -2 \\ f '''(x) = \frac{6}{\left( 1-x \right) ^{4}} \implies f '''(2) = 6 \\ ... \\ \\ f^{(n)}(x) = \frac{n!}{\left( 1-x \right) ^{n+1}} \implies f^{(n)}(2) = (-1)^{n+1}~ n! $$

OpenStudy (owen3):

I assumed \( n = 0,1,2,3... \)

OpenStudy (sh3lsh):

Yep, I think I have this above, and solved it above. Thanks for your help

OpenStudy (sh3lsh):

Mistakenly wrote a in place of 1